# Learned Contextual LiDAR Informed Visual Search in Unseen Environments

### v5

@@ Abstract
This paper presents LIVES: LiDAR Informed Visual Search, an autonomous planner for target search in unknown environments. We consider the pixel-wise environment perception problem where one is given wide Field of View 2D scan data and must perform LiDAR segmentation to contextually label points in the surroundings. These pixel classifications provide an informed prior on which to plan next best viewpoints during visual search tasks. The map-generalizable classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier. An autonomous exploration planner takes the contextual data from scans and uses that prior to plan viewpoints more likely to yield detection of the search target. In order to achieve this, we propose a utility function that accounts for traditional metrics like information gain and path cost and also for the additional contextual information from the scan classifier. LIVES is baselined against several existing exploration methods in simulation to verify its performance. Finally, it is validated in real-world experiments searching for single and multiple targets with a Spot robot in two unseen environments. Videos of experimental validation, implementation details and open source code can be found on our project website at https://sites.google.com/view/lives-2024/home.
@@ I INTRODUCTION
Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [35], target search [1, 12], disaster response [15, 40, 4], agricultural monitoring [13], reconnaissance and surveillance [16, 20, 23]. Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments. This requires planners equipped to generate high quality plans under uncertainty. Conventional methods for LiDAR and multi sensor exploration effectively map environments with focus on completeness, rather than being tailored time critical missions, which we investigate in this paper.
@@
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. Use of scans is often limited to obstacle avoidance, localization and mapping. In contrast, state of the art exploration planners like [3, 42] demonstrate clear improvements by fusing LiDAR and visual information. At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing. While several methods exist for 3D LiDAR scan segmentation [29, 30, 45], these methods are limited by the high computational requirements to handle 3D cloud data [17]. Real time projection-based methods instead aim to detect moving obstacles [11] or objects to be avoided in 3D scans, often for autonomous driving applications [21, 2, 26]. There are also numerous planners designed to reduce uncertainty in new environments by exploiting features common to the environments in which they are designed to operate [33, 31]. Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [3, 14, 36, 42], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [43]. There exists a gap of visual search planners utilizing additional information available in LiDAR scans from segmentation. This work combines these concepts to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
@@
Figure 1: Overview of a planning step of LIVES. In (a) raw 2D LiDAR scans, pose information and image data are received from the Spot robot. They are sent through a neural network classifier III-C (b) to label each point as map or non-map (c). Separately, LiDAR scans and pose data are used to maintain the global search maps (d). Global search maps are used to generate frontier points (e), which are used for finding candidate viewpoints. Finally, each frontier point is sampled at four viewpoints (f) and scored III-E2. In this toy example the left viewpoint in (f) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed. Finally, the selected viewpoint is sent back to the robot for execution.
@@
The key insight in this work is to exploit contextual information available in wide Field of View (FoV) LiDAR scans to filter out points unlikely to be the search target. In particular, this method deviates from traditional exploration methods, which focus on completeness. Rather, LIVES prioritizes regions more likely to be the search target in the robot’s surroundings. We hypothesize exploring agents can improve time efficiency in visual search by exploiting this additional information. We accomplish this using a map based classifier to train a model to recognize features inherent to many structured, indoor environments. The perception model feeds the planner those points unlikely to be permanent, non-map points, for use during viewpoint planning. These points guide the NBV planner towards possible search target locations, assumed to be in the set of non-map points in the environment. Priority is given to features of interest using a novel utility function that accounts for both traditional metrics and non-map points. This information obviates the need to visually inspect parts of the environment that are likely to be permanent. An overview of the proposed method can be found in Fig. 1.
@@
  • Deploy and verify the proposed method on a real robotic system in multiple unseen indoor environments
@@ II-A Next-Best View and Multisensor Planning
Exploration is often solved using next-best viewpoint planners like Frontier Exploration [44, 14]. A common approach uses an active perception formulation to maximize entropy reduction [19, 10]. Ref. [12] combine frontier and sampling based methods to reduce computational overheard for deployment on limited compute UAVs. Ref. [9] learns a policy to generate non-myopic plans online by estimating potential gains of exploring different areas. Ref. [22] similarly learn an information aware policy offline that proposes waypoints for the robot. These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information. Ref. [46] propose a POMDP planner for generalized local object search in cluttered 3D environments and extend it with a 2D planner for larger areas. In contrast to LIVES, their work is aimed at 3D search for small objects, rather than quickly searching buildings during a search and rescue scenario.
Multisensor exploration involves the inclusion of LiDAR information in planning for visual exploration tasks. State of the art multisensor exploration with multiple UAVs [3] leverages the single agent volumetric exploration algorithm presented in [42]. These methods combine frontier generation for both sensors and use a single utility function to score frontier candidates for surface mapping. Ref. [27] maintains a dense 3D pointcloud for recognizing of objects in subterranean tunnel environments. These classifications are utilize by a Pan-Tilt-Zoom camera to perform object detection in Search and Rescue applications. These multisensor approaches demonstrate efficient results in a number of real world environments, however none of these methods is tailored for the indoor environments considered in this article.
@@ II-B World Prediction Based Planning
Often world prediction methods involve learning a model for relevant environments and leveraging their structure for exploring high-value regions first. Ref. [33] use topological features from a database of worlds to inform a frontier-based exploration policy and increase the rate of area explored. Prediction involves recognizing features common to tunnel networks to improve exploration strategies. Similarly, [36] uses a prediction module that detects semantic objects, classifies frontiers, and predicts information gain beyond those frontiers. The planner uses this to explore rooms fully before continuing global exploration. This contrasts from the proposed method that aims to ignore low value regions in order to reduce search completion time. Ref. [31] instead consider visual navigation with occupancy anticipation from egocentric RGB-D based visual context. Ref. [8] learn to predict locations of exits in building with a convolutional neural network (CNN) using a database of building blueprints. Prediction guide the robot towards likely exit locations.
@@ II-C Lidar Classification
Indoor pointcloud classification is tackled in 3D [39] using the KPConv architecture [37]. This method requires the offline annotation and map building at instantiation, and is used for localization and navigation in later work [38]. Neither work validates on a real robot platform. The proposed work instead focus on low computation 2D scan pixel classification.Ref. [28] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [39], focused on map building rather than contextual labeling for planning. Finally, [34] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar points, which are used in a feature extraction process for mapping outdoor environments. In contrast LIVES recognizes non-map features for indoor environments.
@@ III Methods
The overall scan classifier and planning module are described in Alg. 1. The key components are detailed in this section.
@@ III-B Ground Truth LiDAR Scan Classification
The search target of interest belongs to $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$. In particular, scan classification attempts to divide $\mathcal{E}^{\text{occ}}$ into two subsets, namely $\mathcal{E}^{\text{non-map}}$ and $\mathcal{E}^{\text{map}}$. In this work, $\mathcal{E}^{\text{map}}$ are points in the environment that are deemed to be Long-Term Features (LTFs) during classification, while $\mathcal{E}^{\text{non-map}}$ are points that are either Short-Term Features (STFs) or Dynamic Features (DFs) [7]. LTFs represent the static map obstacles, STFs represent static un-mapped obstacles and DFs represent moving points. Given map information, scan points can be classified into these three categories as follows.
@@ III-B1 LTF
First, an analytic ray cast is performed [6] to determine expected laserscan based on map $\mathcal{M}$ and current robot position $x_{i}$. Given observations, the probability that points correspond to one of the lines of that static map can be written
@@ III-C1 Dataset and Data Acquisition
The collected dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model. The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan. $N$ is the total number of data points collected. Collecting a dataset using a real robot requires a human expert to control the robot over long periods of time. This approach is difficult to scale due to mechanical limits of the robot and difficulty of control over large, potentially out-of-sight places. To ease data acquisition we design a steerable cart platform (see website for details) to mimic the robot. Human operators easily maneuver this cart to generate the dataset without strain to the robot. We mount a Velodyne Puck LiDAR for recording scan information and a front-mounted Intel RealSense T265 stereo camera to obtain the cart’s odometry estimate, required for running the ground-truth classification. Finally, a laptop is used to maintain the pose estimate [7], collecting the data and generating online ground-truth labels for the dataset. This laptop records 1) robot pose at each timestep, 2) raw scan data, and 3) classified scan from the ground-truth module for dataset $\mathcal{D}$. The Puck has $n_{i}=897$ and classified scans are recorded at 5Hz. A dataset of size $N\simeq 145,000$ is collected in this study, representing roughly 8 hours of time. In practice, the model only receives raw scan data, pose estimate and previously predicted labels from the robot during operation and must classify pixel-wise the scan online. Training data can be found on the project website.
@@ III-C2 Architecture and Training
We formulate pixel-wise LiDAR classification as a supervised learning problem during training and an auto-regressive problem during inference. Given dataset $\mathcal{D}$, we shuffle and batch the dataset between time steps and locations to minimize location-based and temporal bias during training. Because no map information is provided explicitly, we provide the model with a history buffer of $k$ time steps for scan ranges and robot poses, and $k-1$ previously estimated ground truth labels. Because we do not have access to ground truth labels at inference time, we corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution. We then take the exponential weighted average of the last $k-1$ corrupted pixel-wise classifications to estimate the current pixel-wise classification. This estimate is concatenated with the $k-1$ corrupted ground truth labels to produce a $k$-length matrix of estimated scan labels. During inference, the label history buffer is populated by previous estimates from the policy, cropped by a length $k$ sliding-window. The history buffer is then concatenated with the updated pixel-wise exponential weighted average. At the initialization of inference, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is initialized. In practice there are a large number of non-map points due to the size of scans ($n_{i}=897$), therefore we apply filtering steps like pooling and removing those in previously inspected areas before these points are used for planning. Code for training a model can be found on the project website.
Figure 2: Pixel-wise LiDAR scan classification model architecture. $k$ is the length of the history buffer. (a) $[3,k]$ pose history matrix containing $[x,y,\theta]$. (b) $[n_{i},k]$ LiDAR range history matrix. (c) $[n_{i},k-1]\cup[n_{i},1]$ estimated label history matrix concatenated with its pixel-wise exponential weighted average. (d) The model consists of three temporal-convolutional encoders (TCN), one for each input. The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan. In (e) a threshold is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
@@ III-D Map Updates
Two global occupancy maps (search maps) are maintained. The first is a LiDAR global search map and the second is a visual global search map. Global search maps are implemented in this work using ROS Costmap2D [24]. All cells in both global maps are initialized to be unknown, $e=-1$ $\forall e\in\mathcal{E}$. The robot is equipped with a visual sensor with a cone shaped FoV, implemented as a triangular costmap. The LiDAR sensor has a 360 degree FoV, represented as a square-type costmap. Observations for both sensors are a set of range estimates to the nearest occupied point $e\in\mathcal{E}^{\text{occ}}$. Points between the robot and the nearest occupied point along each ray of the LiDAR scan are free points $e\in\mathcal{E}^{\text{free}}$. Given incoming observations from the robot, each global costmap is updated as free space $e=0$ or occupied $e=1$. This process is shown in Fig. 1(d).
@@ III-E1 Candidate Viewpoint Generation
Frontier points in each of the two global search maps are generated using a process similar to [44], shown in Fig. 1(e). Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers. Any group of frontier points beyond a minimum size becomes its own frontier. Frontier centroids of each frontier in both the visual and LiDAR search maps are found by clustering. Four candidate viewpoints are evaluated at each frontier centroid to account for the limited FoV of the vision sensor.

### v8

# Fast LiDAR Informed Visual Search in Unseen Indoor Environments

@@ Abstract
This paper explores the problem of planning for visual search without prior map information. We leverage the pixel-wise environment perception problem where one is given wide Field of View 2D scan data and must perform LiDAR segmentation to contextually label points in the surroundings. These pixel classifications provide an informed prior on which to plan next best viewpoints during visual search tasks. We present LIVES: LiDAR Informed Visual Search, a method aimed at finding objects of interest in unknown indoor environments. A robust map-free classifier is trained from expert data collected using a simple cart platform equipped with a map-based classifier. An autonomous exploration planner takes the contextual data from scans and uses that prior to plan viewpoints more likely to yield detection of the search target. We propose a utility function that accounts for traditional metrics like information gain and path cost and for the contextual information. LIVES is baselined against several existing exploration methods in simulation to verify its performance. It is validated in real-world experiments with single and multiple search objects with a Spot robot in two unseen environments. Videos of experiments, implementation details and open source code can be found at https://sites.google.com/view/lives-2024/home.
@@ I INTRODUCTION
Planning and real-world execution for autonomous visual exploration are receiving significant attention from the robotics community due to their relevance in a wide range of scenarios including inspection [tan2021comprehensive], target search [balta2017integrated, chen2023hybrid], disaster response [delmerico2017active, tomic2012toward, bi2023cure], agricultural monitoring [christiansen2017designing], reconnaissance and surveillance [grocholsky2006cooperative, liang2022reconnaissance, maini2020visibility]. Active planners in unknown spaces are required to make decisions based on incomplete and noisy information about partial environments. This requires planners equipped to generate high quality plans under uncertainty.
@@
Today’s robots are frequently equipped with LiDAR sensors that cast a full view of the surroundings. State of the art exploration planners like [best2022resilient, vidal2020multisensor] demonstrate efficiency by fusing LiDAR and visual information. At the same time, the potential to use additional information from LiDAR scans for online control has grown with advances in parallel computing. While several methods exist for 3D LiDAR scan segmentation [qi2017pointnet, qi2017pointnet++, zamorski2020adversarial], these methods are limited by high computational requirements [jhaldiyal2023semantic]. In contrast, real time projection-based methods often detect moving obstacles to be avoided for autonomous driving applications [chen2022automatic, liu2022less, baur2021slim, mersch2022receding]. Numerous planners are designed to reduce uncertainty in new environments by exploiting features common to places they are designed for [saroya2020online, ramakrishnan2020occupancy].
Figure 1: LIVES allows the Spot robot to search for objects quickly in varied indoor environments. Video of experiments can be found at our website.
Despite improvements in efficiency of planning algorithms for both LiDAR and vision sensors [best2022resilient, dai2020fast, tao2023seer, vidal2020multisensor], it remains an open problem to enhance performance in autonomous planning and execution for the task of visual search [wang2019autonomous]. This work aims to reduce uncertainty in indoor environments by segmenting 2D LiDAR scans online for utilization by a next-best view (NBV) planner.
@@
Figure 2: Overview of a planning step of LIVES. In (a) raw 2D LiDAR scans, pose information and image data are received from the Spot robot. They are sent through a neural network classifier III-C (b) to label each point as map or non-map (c). Separately, LiDAR scans and pose data are used to maintain the global search maps for the two sensors (d). Global search maps are used to generate frontier points, shown in cyan and pink. Frontier centroids are shown as white dots and used for sampling. Each centroid has four viewpoints scored (e) III-E2. In this toy example the left viewpoint in the lower figure in (e) will be scored highest and selected as the next waypoint due to the high number of unknown cells and non-map points observed. Finally, the selected viewpoint is sent back to the robot for execution.
@@
The key insight in this work is to exploit contextual information available in wide Field of View (FoV) LiDAR scans to filter out points unlikely to be the search target. This method deviates from traditional exploration methods, which focus on completeness. Rather, LIVES prioritizes regions more likely to be the search target in the robot’s surroundings. We hypothesize exploring agents can improve time efficiency in visual search by exploiting this additional information. A map based classifier is used to train a model to recognize features inherent to many indoor environments. The perception model feeds the planner the non-permanent points, non-map points, for use during planning. These points guide the NBV planner towards possible search target locations, assumed to be in the set of non-map points in the environment. Priority is given to features of interest using a novel utility function that accounts for both traditional metrics and non-map points. Examples of search environments are shown in Fig. 1. An overview of the proposed method can be found in Fig. 2.
@@ II-A Next-Best View and Multisensor Planning
Exploration is often solved using next-best viewpoint planners like Frontier Exploration [yamauchi1997frontier, dai2020fast]. A common approach uses an active perception formulation to maximize entropy reduction [julian2014mutual, charrow2015information]. Ref. [chen2023hybrid] combine frontier and sampling based methods while other methods leverage learning to plan [cao2023ariadne, lodel2022look]. These methods demonstrate efficient exploration in simulation and on hardware, however none are explicitly designed for visual search using contextual information. Ref. [zheng2023system] propose a POMDP for local object search in cluttered 3D environments and extend it with a 2D planner for larger areas, however, their work is aimed at 3D search for small objects, rather than quickly searching large areas.
Multisensor exploration includes LiDAR information in planning for visual exploration tasks. State of the art multisensor exploration with multiple UAVs [best2022resilient] leverages the single agent exploration algorithm presented in [vidal2020multisensor]. These methods combine frontier generation for both sensors and use a single utility function to score frontier candidates for surface mapping. Ref. [patel2022lidar] maintains a dense 3D pointcloud for recognizing of objects in subterranean tunnel environments, which are used by a Pan-Tilt-Zoom camera. These works demonstrate efficiency in a number of real world environments, however none of these methods is designed for the environments considered in this article.
@@ II-B World Prediction Based Planning
World prediction methods typically learn a model of relevant environments and exploit their structure to explore high-value regions first. Ref. [saroya2020online] use topological features from a database of tunnel networks to inform a frontier-based exploration policy. Similarly, ref. [tao2023seer] use a prediction module to detect semantic objects and classify frontiers, resulting in exploring rooms first before continuing global exploration. The proposed method instead aims to ignore low value regions in order to reduce search completion time. Prediction guide the robot towards likely exit locations.
@@ II-C LiDAR Classification
Indoor 3D pointcloud classification is tackled in [thomas2021self] with KPConv [thomas2019kpconv]. This requires the offline annotation and map building at instantiation and is used for simulated localization and navigation [thomas2022learning]. The proposed work instead focus on low computation 2D scan pixel classification.Ref. [pomerleau2013comparing] presents an in depth comparison and framework for Iterative Closest Point (ICP) methods, similar to that of [thomas2021self], focused on map building. Finally, [shan2018lego] present a map-free framework for segmenting 3D pointclouds into ground, edges and planar pointsfor mapping outdoor environments.
@@ III Methods
The overall scan classifier and planning module are described in Fig. 2 and Alg. 1. The key components are detailed in this section.
@@ III-B Ground Truth LiDAR Scan Classification
The search target of interest belongs to $\mathcal{E}^{\text{occ}}\subset\mathcal{E}$. In particular, scan classification attempts to divide $\mathcal{E}^{\text{occ}}$ into two subsets, namely $\mathcal{E}^{\text{non-map}}$ and $\mathcal{E}^{\text{map}}$. In this work, $\mathcal{E}^{\text{map}}$ are points in the environment that are deemed to be Long-Term Features (LTFs) during classification, while $\mathcal{E}^{\text{non-map}}$ are points that are either Short-Term Features (STFs) or Dynamic Features (DFs) [biswas2017episodic]. LTFs represent the static map obstacles, STFs represent static un-mapped obstacles and DFs represent moving points. Given map information, scan points can be classified into these three categories as follows.
@@ III-B1 LTF
First, an analytic ray cast is performed [biswas2012depth] to determine expected laserscan based on map $\mathcal{M}$ and current robot position $x_{i}$. Given observations, the probability that points correspond to one of the lines of that static map can be written
@@ III-C1 Dataset and Data Acquisition
The dataset $\mathcal{D}$ consists of tuples of data from the ground truth classification model. The tuples are given as $\mathcal{D}=\{(x_{i},s_{i},s_{i}^{\text{class}})\}^{N}_{i=1}$, where $x_{i}$ represents robot pose, $s_{i}$ is the raw LiDAR scan data, and $s_{i}^{\text{class}}$ is the classified scan. $N$ is the total number of data points collected. Data acquisition with a robot requires a human expert to control, which is difficult to scale due to mechanical limits of the robot. To ease data acquisition we design a steerable cart platform (see website for details) to mimic the robot. Human operators easily maneuver this cart to obtain the dataset. We mount a Velodyne Puck LiDAR for recording scan information and a front-mounted Intel RealSense T265 stereo camera to obtain the cart’s odometry estimate, required for running the ground-truth classification. A laptop maintains the pose estimate [biswas2017episodic] and collects the data. The Puck has $n_{i}=897$ and classified scans are recorded at 5Hz. A dataset of size $N\simeq 145,000$ is collected in this study, representing roughly 8 hours of time. In practice, the model only receives raw scan data, pose estimate and previously predicted labels from the robot during operation and must classify pixel-wise the scan online. Training data can be found on the project website.
@@ III-C2 Architecture and Training
We formulate pixel-wise LiDAR classification as a supervised learning problem during training and an auto-regressive problem during inference. Given dataset $\mathcal{D}$, we shuffle and batch the dataset between time steps and locations to minimize location-based and temporal bias during training. Because no map information is provided explicitly, we provide the model with a history buffer of $k$ time steps for scan ranges and robot poses, and $k-1$ previously estimated ground truth labels. Because we do not have access to ground truth labels at inference time, we corrupt the ground truth labels during training by randomly bit-flipping the classification of 10% of the labels, chosen by sampling indices from a uniform random distribution. We then take the exponential weighted average of the last $k-1$ corrupted pixel-wise classifications to estimate the current pixel-wise classification. This estimate is concatenated with the $k-1$ corrupted ground truth labels to produce a $k$-length matrix of estimated scan labels. During inference, the label history buffer is populated by previous estimates from the policy. The history buffer is concatenated with the updated pixel-wise exponential weighted average. At initialization, the policy is bootstrapped with zero-value poses, scans, and labels, and run for $k$ steps of inference until the history buffer is full. Code for training a model can be found on the project website.
Figure 3: Pixel-wise LiDAR scan classification model architecture. This model enables speeding up search by providing contextual information to the planner. $k$ is the length of the history buffer. (a) $[3,k]$ pose history matrix containing $[x,y,\theta]$. (b) $[n_{i},k]$ LiDAR range history matrix. (c) $[n_{i},k-1]\cup[n_{i},1]$ estimated label history matrix concatenated with its pixel-wise exponential weighted average. (d) The model consists of three temporal-convolutional encoders (TCN), one for each input. The encoded poses, scans, and labels are combined to produce a pixel-wise classification of the LiDAR scan. In (e) a threshold (positive/negative) is applied to the raw logits such that each pixel is classified as either a map point or a non-map point.
@@ III-D Map Updates
Two global occupancy maps (search maps) are maintained. One for the LiDAR and the other for the visual sensor. Global search maps are implemented using ROS Costmap2D [costmap]. All cells in both global maps are initialized to be unknown, $e=-1$ $\forall e\in\mathcal{E}$. The robot is equipped with a visual sensor with a cone shaped FoV, implemented as a triangular costmap. The LiDAR sensor has a 360 degree FoV, represented as a square-type costmap. Observations for both sensors are a set of range estimates to the nearest occupied point $e\in\mathcal{E}^{\text{occ}}$. Points between the robot and the nearest occupied point along each ray of the LiDAR scan are free points $e\in\mathcal{E}^{\text{free}}$. Given incoming observations from the robot, each global costmap is updated as free space $e=0$ or occupied $e=1$. This process is shown in Fig. 2(d).
@@ III-E1 Candidate Viewpoint Generation
Frontier points in each of the two global search maps are generated using a process similar to [yamauchi1997frontier], shown in Fig. 2(e). Any free cell ($e=0$) adjacent to an unknown cell ($e=-1$) is a frontier edge, which are grouped together into frontiers. Any group of frontier points beyond a minimum size becomes its own frontier. Frontier centroids of each frontier in both the visual and LiDAR search maps are found by clustering. Four candidate viewpoints are evaluated at each frontier centroid to account for the limited FoV of the vision sensor.